Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability Strategy

Expert replies
by myfish » Fri Apr 27, 2012 10:09 am
There are exactly two ways to attack most probability questions. A) You calculate with P(A) = 1 - P(nA) or B) P(A) directly.

The following example will show what I mean.

A fair six-sided dice with faces numbered one through six rolled three times. What is the probability that the face with the number 6 on it will NOT be facing upward on all three rolls?

Solution A)
P(A) = 1 - P ---> P(6) = 1/6 x 3 rolls = 1/216 ---> AGAIN P(A) = 1 - P ---> 1 - 1/216 = 215/216

Solution B)
P(n6) = 5/6 x 3 rolls = 125/216

Since only one solution is correct and I repeatedly stumble over this very easy issue, I would like to ask the experts how to identify when to use which solution. Anyone any idea? It bugs me to screw up on such easy problems. Thanks so much!
Join the discussion
Source: — Problem Solving |

by Stuart@KaplanGMAT » Fri Apr 27, 2012 10:50 am
myfish wrote:There are exactly two ways to attack most probability questions. A) You calculate with P(A) = 1 - P(nA) or B) P(A) directly.

The following example will show what I mean.

A fair six-sided dice with faces numbered one through six rolled three times. What is the probability that the face with the number 6 on it will NOT be facing upward on all three rolls?

Solution A)
P(A) = 1 - P ---> P(6) = 1/6 x 3 rolls = 1/216 ---> AGAIN P(A) = 1 - P ---> 1 - 1/216 = 215/216

Solution B)
P(n6) = 5/6 x 3 rolls = 125/216

Since only one solution is correct and I repeatedly stumble over this very easy issue, I would like to ask the experts how to identify when to use which solution. Anyone any idea? It bugs me to screw up on such easy problems. Thanks so much!
Hi!

Both solutions are always correct - if you get two different answers, then you must have made a mistake along the way.

Here, for example, you've miscalculated what you call solution B.

In your solution, you only look at the case in which none of them is a 6; however, we also satisfy the question's requirements if 1 or 2 of the dice is a 6.

To properly use solution B, you need to look at the different cases that match your desired results:

0 6s (which you calculated);
1 6; or
2 6s.

Since we're happy with any of those 3 cases, we need to ADD their individual probabilities to get the total probability of "not all 3 6s".

The probability of getting 1 six on 3 dice is:

(1/6)(5/6)(5/6) * 3C1 (we multiply by 3C1 since there are 3 different ways to get 1 six on three dice) = 25/216 * 3 = 75/216

The probability of getting 2 sixes on 3 dice is:

(1/6)(1/6)(5/6) * 3C2 = 5/216 * 3 = 15/216

and finally we add the 3 cases together to get:

125/216 + 75/216 + 15/216 = 215/216.

* * *

As you can see, the "one minus" approach was much quicker. When deciding with of the two approaches to use, just ask yourself "which one involves fewer and/or easier calculations?"
Image

Stuart Kovinsky | Kaplan GMAT Faculty | Toronto

Kaplan Exclusive: The Official Test Day Experience | Ready to Take a Free Practice Test? | Kaplan/Beat the GMAT Member Discount
BTG100 for $100 off a full course
Join the discussion

by myfish » Fri Apr 27, 2012 11:07 am
Thanks so much Stuart.

[quote="Stuart Kovinsky"][quote="myfish"]There are exactly two ways to attack most probability questions. A) You calculate with P(A) = 1 - P(nA) or B) P(A) directly.

The following example will show what I mean.

A fair six-sided dice with faces numbered one through six rolled three times. What is the probability that the face with the number 6 on it will NOT be facing upward on all three rolls?

Solution A)
P(A) = 1 - P ---> P(6) = 1/6 x 3 rolls = 1/216 ---> AGAIN P(A) = 1 - P ---> 1 - 1/216 = 215/216

Solution B)
P(n6) = 5/6 x 3 rolls = 125/216

Since only one solution is correct and I repeatedly stumble over this very easy issue, I would like to ask the experts how to identify when to use which solution. Anyone any idea? It bugs me to screw up on such easy problems. Thanks so much![/quote]

Hi!

Both solutions are always correct - if you get two different answers, then you must have made a mistake along the way.

Here, for example, you've miscalculated what you call solution B.

In your solution, you only look at the case in which none of them is a 6; however, we also satisfy the question's requirements if 1 or 2 of the dice is a 6.

To properly use solution B, you need to look at the different cases that match your desired results:

0 6s (which you calculated);
1 6; or
2 6s.

Since we're happy with any of those 3 cases, we need to ADD their individual probabilities to get the total probability of "not all 3 6s".

The probability of getting 1 six on 3 dice is:

(1/6)(5/6)(5/6) * 3C1 (we multiply by 3C1 since there are 3 different ways to get 1 six on three dice) = 25/216 * 3 = 75/216

The probability of getting 2 sixes on 3 dice is:

(1/6)(1/6)(5/6) * 3C2 = 5/216 * 3 = 15/216

and finally we add the 3 cases together to get:

125/216 + 75/216 + 15/216 = 215/216.

* * *

As you can see, the "one minus" approach was much quicker. When deciding with of the two approaches to use, just ask yourself "which one involves fewer and/or easier calculations?"[/quote]
Join the discussion